# Identifying the Desired Word Suggestion in Simultaneous Audio

### v1

@@ 4.3. Discussion and Limitations
While the results shown in these studies (particularly the latter) are promising, they are still only a simplification of a real-world text entry task. The studies presented here allowed participants to focus all of their attention on listening to the suggestions for the target word. In a more realistic task, users attention will be split between not only the word suggestions, but also the content that they are writing and the act of performing the input itself. This will increase the cognitive load, and we suspect this may further decrease user performance.
An additional limitation is that in these studies, participants were instructed to wear headphones, but we were not able to detect if participants followed this instruction. Additionally, in a real-world task, wearing headphones may not always be practical, which could impair users’ ability to hear separate left and right audio channels.
Our study consisted of a single study session lasting less than an hour. The ability to successfully detect a target word in simultaneous TTS audio may require more time to learn. It would be interesting to conduct a longitudinal study to investigate how user performance changes with practice.
@@ 5. Conclusion
In this work, we explored different ways to present word suggestions to users using only TTS audio feedback. To our knowledge, we provide the first quantitative results showing if users can perform well in the challenging task of detecting a single target word within a set of similar sounding words. In our first study, we found playing two words simultaneously using spatial audio had a lower detection accuracy of 84–85% compared to 96% for sequential playback (though this difference was not significant). Playing three or four words at the same time was significantly worse than sequential playback. In our second study, we compared adding small delays before starting each simultaneous voice. We found by adding a short 0.15 s delay, users detected the target word about equally as well at 84% using two simultaneous voices versus 86% using two fully sequential voices. Playing two words simultaneously made audio feedback 32% faster. Based on our two studies, if simultaneous audio is to be used for short audio feedback events, we suggest using a small start time offset in addition to spatial audio and to limit the number of voices to two.
@@ Acknowledgements.
This work was supported by NSF IIS-1909248 and by an NSF Graduate Research Fellowship (2034833).

### v2

@@ 4.3. Discussion
4.4. Limitations
While the results shown in our studies (particularly Study 2) are promising, they are still only a simplification of real-world text entry. Our studies allowed participants to focus all of their attention on listening to the suggestions for the target word. In a more realistic task, users attention will be split between not only the word suggestions, but also the content that they are writing and the act of performing the input itself. This will increase the cognitive load, and we suspect this may further decrease user performance.
An additional limitation is that while we instructed participants to wear headphones, we do not know for sure if participants followed this instruction. Additionally, in a real-world task, wearing headphones may not always be practical, which could impair users’ ability to hear separate left and right audio channels.
Our study consisted of a single study session lasting less than an hour. The ability to successfully detect a target word in simultaneous TTS audio may require more time to learn. It would be interesting to conduct a longitudinal study to investigate how user performance changes with practice. We also did not specifically recruit blind or low vision participants, who likely have more experience listening to TTS audio. Additional follow-up work is required to determine how well these results generalize to blind and low vision users.
@@ 5. Conclusion
We explored different ways to present word suggestions to users using only TTS audio feedback. To our knowledge, we provide the first quantitative results showing if users can perform well in the challenging task of detecting a single target word within a set of similar sounding words. In our first study, we found playing two words simultaneously using spatial audio had a lower detection accuracy of 84–85% compared to 96% for sequential playback (though this difference was not significant). Playing three or four words at the same time was significantly worse than sequential playback.
In our second study, we compared adding small delays before starting each simultaneous voice. We found that by adding a short 0.15 s delay, users detected the target word about equally as well at 84% using two simultaneous voices versus 86% using two fully sequential voices. Playing two words simultaneously made audio feedback 32% faster. Based on our two studies, if simultaneous audio is to be used for short audio feedback events, we suggest using spatial audio, a small start time offset, and to limit the number of voices to two.
@@ Acknowledgements.
This work was supported by NSF IIS-1909248 and by NSF Graduate Research Fellowship DGE-2034833. We thank Ricardo Gonzalez for pilot testing the web application and suggesting the use of a delay in the speech.